# In-Context Reverse Classification Accuracy: Efficient Estimation of Segmentation Quality without Ground-Truth

### v1

@@ 3 Experiments and results
Impact of dataset size and RA on performance. We evaluate the impact of the reference dataset size and the use of RA by applying In-Context RCA with UniverSeg as our reverse classifier, considering different support set sizes. We aim to determine if RA maintains or even improves prediction quality while using a smaller reference dataset, thus reducing inference time in practice. We considered reference dataset sizes of 2, 5, 10 and 25 examples. Beyond these values, inference time per image increases substantially. Since UniverSeg supports only binary segmentation, for multi-class we process them separately (including background), and then select the most likely class per pixel using softmax, as described in [4].
@@
Fig. 3 shows the results on HC18 and JSRT datasets, comparing the performance of In-Context RCA for different reference datasets of size $k$. The figure contrasts two selection methods: random selection (first row) and RA (second row), where the $k$ most similar images are chosen based on cosine similarity in the DINOv2 embedding space. The horizontal axis represents the predicted DSC, while the vertical axis shows the real value. As the number of reference samples increases, both methods show improved performance, yielding higher correlation and lower mean absolute error (MAE). However, RA achieves better results with fewer reference samples, what is particularly beneficial during deployment, as it reduces inference time. Other similarity metrics, such as Euclidean distance and inner-product were also considered, but cosine similarity showed the best results.
Figure 3: Scatter plots comparing predicted vs real DSC on HC18 (Left) and JSRT (Right) for different reference datasets of size $k$. Top: results using randomly selected reference samples. Bottom: results following retrieval-augmentation, where the $k$ most similar images are selected based on cosine similarity in the DINOv2 embedding space. In both cases retrieval-augmentation achieves better performance with fewer samples, as shown by higher correlation and lower MAE (bold indicates the best value).
3.1 Comprehensive evaluation across datasets
@@
In order to assess the effectiveness and limitations of In-Context RCA, we conducted an extensive evaluation using UniverSeg and SAM 2 as reverse classifiers, considering DSC as evaluation metric. We compared the performance of In-Context RCA with a single-atlas approach, known to offer the strongest results in the classical RCA framework [31, 24]. For the atlas, we included both traditional and RA versions, whereas for UniverSeg and SAM 2, we focused exclusively on the RA version as it showed better pefromance in our previous experiment (see Fig. 3). We used the segmentation masks generated by a UNet trained for different epochs as previously discussed, to ensure evaluation across varying segmentation qualities. We observe that following RA reference sample selection yields good results without requiring a large reference database.All experiments were conducted on an AMD Ryzen 7 4800H CPU (8 cores) and an NVIDIA GTX 1660 Ti GPU (6GB of VRAM). In all cases, we fixed the size of the reference dataset to 8, based on previous observations.
@@
Fig. 4 provides a comprehensive overview of the results when predicting the DSC across all datasets for all four variants: the traditional RCA framework using single-atlas, its RA version, and the In-Context RCA methods using UniverSeg and SAM 2 as reverse classifiers. We observe that In-Context RCA demonstrates similar and, in some cases, better results compared to the Atlas method, while being significantly more computationally efficient. Particularly in highly irregular segmentation problems like the NuCLS histopathology dataset, our SAM2 variant achieves good correlation of about 0.74 while Atlas methods completely fail. Importantly, while the classic atlas method takes on average one minute to process each image, SAM 2 and UniverSeg achieve the same task in just 0.7 seconds and 0.37 seconds respectively, approximately 85 and 162 times faster. This efficiency, combined with its strong performance, makes In-Context RCA an ideal choice for integration into clinical pipelines.
4 Conclusion
@@
In this work, we addressed the critical challenge of evaluating medical image segmentation quality without available GT. To do so we introduced In-Context RCA, a novel framework that builds on top of RCA, leveraging the power of in-context segmentation models as reverse classifiers. By also incorporating retrieval augmentation, our method enables efficient quality estimation using smaller reference databases while either maintaining or improving prediction accuracy.
@@
Experimental results conducted across diverse medical imaging modalities, show that In-Context RCA achieves comparable or superior performance to traditional atlas-based approaches while being significantly more computationally efficient. It also overcomes the limitations imposed by atlas-based methods allowing to tackle scenarios beyond anatomical segmentation, such as cell images. Overall, our results position In-Context RCA as a promising solution for automated quality control in clinical workflows, where rapid assessment is essential.

### v3

@@ V-B Evaluation of In-Context RCA
Impact of dataset size and RA on performance. We evaluate the impact of the reference dataset size and the use of RA (retrieval augmentation) by applying In-Context RCA with UniverSeg as our reverse classifier, considering different support set sizes. We aim to determine if RA maintains or even improves prediction quality while using a smaller reference dataset, thus reducing inference time in practice. We considered reference dataset sizes of 2, 5, 10 and 25 examples. Larger sets noticeably increase inference time per image without any meaningful improvement in prediction accuracy. Since UniverSeg supports only binary segmentation, for multi-class we process them separately (including background), and then select the most likely class per pixel using softmax, as described in [9].
@@
Fig. 3 shows the results on 3D-IRCAdB and JSRT datasets, comparing the performance of In-Context RCA for different reference datasets of size $k$. The figure contrasts two selection methods: random selection (first row) and RA (second row), where the $k$ most similar images are chosen based on cosine similarity in the DINOv2 embedding space. The horizontal axis represents the predicted DSC, while the vertical axis shows the real value. As the number of reference samples increases, both methods show improved performance, yielding higher correlation and lower mean absolute error (MAE). However, RA achieves better results with fewer reference samples, what is particularly beneficial during deployment, as it reduces inference time. Other similarity metrics, such as Euclidean distance and inner-product were also considered, but cosine similarity showed the best results.
@@
Comprehensive evaluation across datasets. In order to assess the effectiveness and limitations of In-Context RCA, we conducted an extensive evaluation using UniverSeg and SAM 2 as reverse classifiers, considering DSC as evaluation metric. We compared the performance of In-Context RCA with a single-atlas approach, known to offer the strongest results in the classical RCA framework [5, 19]. For the atlas, we included both traditional and RA versions, whereas for UniverSeg and SAM 2, we focused exclusively on the RA version as it showed better pefromance in our previous experiment (see Fig. 3). We used the segmentation masks generated by a UNet trained for different epochs as previously discussed, to ensure evaluation across varying segmentation qualities. We observe that following RA reference sample selection yields good results without requiring a large reference database. In all cases, we fixed the size of the reference dataset to 8, based on previous observations.
@@
Fig. 4 provides a comprehensive overview of the results when predicting the DSC across all datasets for all four variants: the traditional RCA framework using single-atlas, its RA version, and the In-Context RCA methods using UniverSeg and SAM 2 as reverse classifiers. We observe that In-Context RCA demonstrates similar and, in some cases, better results compared to the Atlas method, while being significantly more computationally efficient. Particularly in highly irregular segmentation problems like the NuCLS histopathology dataset, our SAM2 variant achieves good correlation of about 0.70 while Atlas methods completely fail.
Performance gains for In-Context RCA. Importantly, while the classic RCA-atlas method takes on average one minute to process each image, SAM 2 and UniverSeg achieve the same task in just 0.70 seconds and 0.37 seconds respectively, approximately 85 and 162 times faster. This efficiency, combined with its strong performance, makes In-Context RCA an ideal choice for integration into clinical pipelines.
V-C Evaluation of Conformal RCA
To evaluate our conformal RCA approach, we split each dataset into training (80%), calibration (10%) and test (10%) subsets. The calibration set is used to compute the CP threshold $\hat{q}$ that adapts the prediction intervals, while performance is reported on the held-out test set.
In order to enable direct comparison with [26], we adopt PHiSeg [28]—the top performer in that work—as our backbone segmentation model. Importantly, Conformal-RCA remains model-agnostic, making it applicable to any underlying segmentation network regardless of output format. Segmentation masks of varying quality for both calibration and test sets were generated by training PHiSeg and saving intermediate checkpoints, following the same procedure as before. For PHiSeg, we follow the conformal calibration procedure described in [26]: for each input image, $N=50$ stochastic segmentations are generated, and the DSC is estimated from the softmax outputs [32], yielding a distribution of predicted DSCs per sample. Nonconformity scores are defined as the absolute residuals between the mean predicted and ground truth DSC, scaled by the empirical standard deviation of the distribution. We additionally explore PHiSeg (CQR), a newly proposed variant where nonconformity scores are computed following the empirical CQR method. In this case, empirical quantiles are computed at levels 0.2 and 0.8.
Fig. 5 reports the distribution of prediction interval sizes and empirical coverage levels across all datasets for the evaluated methods. We include different versions of our approach, which differ only in the choice of reverse segmenter: the classical atlas-based method, and our two proposed in-context learning-based alternatives—UniverSeg and SAM 2. For all variants, we used a reference database of size 32, selected following the RA approach previously described. Empirical quantiles were computed at levels $p_{\ell}=0.4$ and $p_{h}=0.95$ and prediction intervals were constructed with a user-specified coverage of 90% ($\alpha=0.1$) as described in Section IV-B.
Panel (a) of Fig. 5 shows the distribution of prediction interval widths across all datasets. Among our Conformal-RCA variants, SAM 2 and Atlas achieve consistently tight intervals, whereas UniverSeg yields somewhat wider intervals. Notably, SAM 2 offers a substantial computational speed‑up over the atlas‑based approach while maintaining similar interval quality. Compared to the PHiSeg baseline, UniverSeg occasionally produces wider intervals; however, both SAM 2 and Atlas consistently attain lower median widths and reduced spread, indicating more robust uncertainty quantification. Importantly, compared to the conformal procedure used in prior work [26], the CQR variant of PHiSeg produces tighter intervals, further highlighting the benefits of following our proposed empirical CQR method within the CP framework.
As shown in panel (b) of Fig. 5, neither our Conformal‑RCA nor the PHiSeg variants consistently reach the target marginal coverage of 90%. This under-coverage phenomenon was also observed in [26] and is likely due to violations of the exchangeability assumption between the test and calibration splits. To further support this hypothesis, we ran controlled experiments in which both calibration and test sets are drawn from the same known distribution. In these controlled settings where the exchangeability assumption holds, the desired nominal coverage level of 90% is achieved (see Appendix).
VI Conclusion
@@
In this work, we addressed the critical challenge of evaluating medical image segmentation quality in the absence of ground truth by presenting two complementary frameworks. First, In-Context RCA leverages the power of in-context segmentation models as reverse classifiers and integrates retrieval augmentation to select relevant reference samples. This combination enables efficient quality estimation using smaller reference databases while either matching or improving prediction accuracy of traditional RCA.
Second, Conformal RCA extends the traditional RCA procedure using split conformal prediction to produce prediction intervals with statistical guarantees, providing a more comprehensive assessment of segmentation quality that moves beyond point estimates. Unlike previous interval-estimation approaches that rely on specialized probabilistic architectures and require calibrated softmax outputs, Conformal RCA is fully model‑agnostic and easily integrates with any segmentation network.
@@
Experimental results conducted across diverse medical imaging modalities, show that In-Context RCA achieves comparable or superior performance to traditional atlas-based approaches while being up to two orders of magnitude faster. It also overcomes the limitations imposed by atlas-based methods allowing to tackle scenarios beyond anatomical segmentation, such as cell images. Conformal RCA demonstrates the ability to produce meaningful prediction intervals that outperform existing conformal prediction approaches for segmentation quality control.
A remaining limitation is the reliance of exchangeability between calibration and test sets. When this assumption is violated (e.g., due to domain shifts in scanner settings or patient populations), coverage can degrade. Developing methods that are robust to distribution shifts is an important direction for future work.
Overall, our results position both In-Context RCA and Conformal RCA as promising and practical solutions for automated quality control in clinical workflows, where rapid and reliable segmentation assessment is essential.
Appendix A: Controlled Experiments on Exchangeability Assumption
To illustrate how exchangeability underpins Conformal RCA’s coverage guarantees, we run a controlled experiment in which both calibration and test Dice scores are drawn from the same underlying distribution. RCA scores are then obtained by adding Gaussian noise to these true scores (clipped to [0,1]). We then apply Conformal RCA with the same settings used in Section V-C: a reference set of size 32 and empirical quantiles at levels $p_{\ell}=0.4$ and $p_{h}=0.95$.
Fig. 6 shows the resulting nonconformity‑score histograms for calibration and test splits, along with the computed threshold $\hat{q}$. Empirical marginal coverage surpasses the desired 90 % target level, confirming that Conformal RCA achieves valid coverage under ideal conditions.
Figure 6: Distribution of nonconformity scores for synthetic calibration and test sets, generated by sampling true Dice scores from a $\mathrm{Beta}(2,2)$ distribution and adding noise drawn from $\mathcal{N}(0,0.1)$ to simulate variability in RCA scores. The vertical dashed line indicates the conformal quantile threshold $\hat{q}$.